# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v2

@@ IV-B High-Dimensional Latent State Estimation
IV-C Real-World Time Series Prediction
In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets commonly used for system identification (see Table I). The prediction RMSE results are summarized in Table I, comparing ETGPSSM against SOTA methods, including neural network-based approaches—DKF [7] and AD-EnKF [50]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37]. Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
As shown in Table I, ETGPSSM demonstrates competitive performance across all datasets. Notably, on the Actuator dataset, ETGPSSM (dnn) and ETGPSSM (bnn) both achieve an improved RMSE of 0.646, outperforming EnVI (0.657) and significantly improving upon AD-EnKFs, which record an RMSE of 0.705. Similarly, on the Ball Beam and Gas Furnace datasets, ETGPSSMs attain the first and second lowest RMSEs, highlighting their ability to accurately model system dynamics. However, on datasets such as Drive and Dryer, other GPSSMs like PRSSM and EnVI achieve slightly superior performance. This can be attributed to the relatively low latent state dimension ($d_{x}\!=\!4$), where traditional GPSSMs, such as PRSSM, ODGPSSM and EnVI, remain highly competitive, as discussed in [37]. Nonetheless, we emphasize that the proposed ETGPSSM remains highly competitive, achieving very comparable performance even in low-dimensional settings, while significantly reducing computational complexity and parameter overhead—thereby ensuring both scalability and efficiency in high-dimensional scenarios.
In comparison to neural network-based methods like AD-EnKF, the superior performance of ETGPSSM is driven by its ability to integrate the strengths of GPs and neural networks. This hybrid approach allows ETGPSSM to effectively model complex, non-stationary system dynamics while providing reliable predictions—an area where pure deep neural network-based methods often struggle. This is evident from the results, where neural network-based methods exhibit higher prediction errors. Notably, AD-EnKF (bnn) generally outperforms other network-based approaches, corroborating the benefits of Bayesian treatments, particularly their inherent regularization, which helps mitigate overfitting and improve generalization. In summary, the experimental results demonstrate that ETGPSSM is a powerful and versatile tool for real-world time series prediction.
@@ V Conclusions
In this paper, we presented a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our key contribution is the introduction of the ETGP prior, which integrates a shared GP with input-dependent normalizing flows modeled via neural networks. This ETGP enhances flexibility in capturing complex, non-stationary transitions while significantly reducing the computational and parametric complexity of traditional GPSSMs in high-dimensional settings. To enable scalable inference, we developed a variational framework that approximates the posterior over the ETGP by separately modeling the underlying GP and neural network parameters. Additionally, we leverage the EnKF to circumvent the need for explicit parameterization of state posteriors, further improving computational efficiency. Empirical evaluations across dynamical system learning, high-dimensional state estimation, and time-series prediction tasks demonstrate that our method outperforms existing GPSSMs and deep neural network-based SSMs, while maintaining lower computational and parametric complexity.
Appendix A Proof of Corollary 1
@@
Appendix B ELBO Derivations and Approximations
Given the joint distribution of our model in Eq. (18) and the variational distribution in Eq. (21), augmented by the sparse GP, we can formulate the ELBO according to the general definition provided in Eq. (17). The detailed derivations are presented in Eq. (45).
From Eq. (45d) to Eq. (45e), we have made the following two assumptions, similar to the work in [37]:
  • 1)
    $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$,
  • 2)
    $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$.
The transition from Eq. (45f) to Eq. (45g) is derived straightforwardly by applying Bayes’ theorem.
The two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP. The rationale behind these approximations is as follows:
$\mathcal{L}$ $=\mathbb{E}_{q}\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}{q({\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}\right]$ (45a)
@@ Appendix B ELBO Derivations and Approximations
$=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45g)

### v3

@@ IV-C Real-World Time Series Prediction
As reported in Table I, ETGPSSM demonstrates strong overall performance. On the Actuator dataset, both ETGPSSM (dnn) and ETGPSSM (bnn) achieve an RMSE of $0.646$, outperforming EnVI ($0.657$) and significantly improving upon AD-EnKFs, which yield an RMSE of $0.705$. Similar trends are observed on the Ball Beam and Gas Furnace datasets, where ETGPSSM ranks first and second, respectively, indicating effective modeling of system dynamics. On the Drive and Dryer datasets, however, GPSSMs such as PRSSM and EnVI attain slightly better performance. This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [37]. Nonetheless, we highlight that although the primary contribution of the proposed ETGPSSM lies in its scalability and efficiency in high-dimensional settings, it also remains highly competitive in low-dimensional scenarios, achieving superior or very comparable performance.
In comparison to neural network-based methods like AD-EnKF, ETGPSSM’s hybrid design leverages the representational power of deep learning while retaining the uncertainty quantification and regularization benefits of GPs. This synergy proves especially effective for modeling complex, non-stationary dynamics. This is evident from the results, where neural network-based methods exhibit higher prediction errors. AD-EnKF (bnn) consistently outperforms its DNN counterpart, underscoring the robustness provided by Bayesian approaches. Still, both lag behind ETGPSSM in predictive accuracy across most datasets. In summary, ETGPSSM demonstrates strong predictive performance on diverse real-world time series, matching or outperforming prior methods, while offering a scalable and efficient solution applicable to both low- and high-dimensional dynamical systems.
@@ V Conclusions
In this paper, we proposed a new framework for learning and inference in GPSSM tailored to non-stationary, high-dimensional dynamical systems. Our central contribution is the introduction of the ETGP prior, which combines a shared GP with input-dependent normalizing flows parameterized by neural networks. This construction enables expressive, non-stationary modeling while reducing the computational and parametric overhead typical of high-dimensional GPSSMs. To support scalable inference, we developed a variational framework that approximates the ETGP posterior by disentangling the GP and neural network components. The generic and assumption-free variational lower bound we derive requires no explicit latent state posterior parameterization, but instead uses EnKF for the inference, yielding additional computational efficiency. Extensive experiments across system identification, high-dimensional filtering, and time series forecasting show that ETGPSSM consistently achieves strong performance while maintaining favorable computational efficiency.
Appendix A Proof of Proposition 1
Proof.
Our target is to compare the two lower bounds on the log-evidence $\log p({\mathbf{y}}_{1:T})$:
$\mathcal{L}$ $=\mathbb{E}_{q({\mathbf{x}}_{0},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0},\bm{f})}\right],$ (41)
$\mathcal{L}_{2}$ $=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})}{q({\mathbf{x}}_{0:T},\bm{f})}\right].$ (42)
Since
$q({\mathbf{x}}_{0:T},\bm{f})=q({\mathbf{x}}_{0},\bm{f})q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0}),$
we have $\mathcal{L}_{2}=$
$\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})p({\mathbf{x}}_{1:T}|{\mathbf{x}}_{0},\bm{f},{\mathbf{y}}_{1:T})}{q({\mathbf{x}}_{0},\bm{f})q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})}\right]$ (43)
$\!=\!\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0})q(\bm{f})}+\log\frac{p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})}{q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})}\right]$
$\!=\!\mathcal{L}-\mathbb{E}_{q(\bm{f},{\mathbf{x}}_{0})}\left[\text{KL}\big{(}q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\|p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})\big{)}\right]$
Since KL divergence is non-negative:
$\text{KL}\big{(}q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\|p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})\big{)}\geq 0$
with equality iff $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})=p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$. Thus we have:
$\mathcal{L}\geq\mathcal{L}_{2}.$
Equality holds when $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})=p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$ for all $\bm{f}$ and ${\mathbf{x}}_{0}$. ∎
Appendix B ELBO Derivation
Proof.
We derive the ELBO $\mathcal{L}$ for our ETGPSSM by first considering a general stochastic process prior $\bm{f}$ over the transition function (Eq. (22)), then specializing to our ETGP construction. The derivation proceeds as follows:
$\mathcal{L}$ $=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})}{q({\mathbf{x}}_{0},\mathbf{w},\tilde{f})}\right],$ (44a)
$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f})}{q({\mathbf{x}}_{0})q(\mathbf{w})q(\tilde{f})}\right]$ (44b)
$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (44c)
$=\mathbb{E}_{q}\left[\log p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
$\quad-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (44d)
$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
$\quad-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (44e)
where we augment the GP using inducing points and use the variational sparse GP in Eq. (44c); We factorize the joint likelihood into sequential terms in Eq. (44e), where for $t=1$ we have the log-term as $\log p({\mathbf{y}}_{1}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})$. ∎
Appendix C Proof of Corollary 1
@@
$\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (50)
Supplement D Discussions about ELBO Derivation
Following the GPSSM literature [37, 31], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
$\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right].$ (51)
It turns out that $\mathcal{L}_{2}$ is a lower bound of $\mathcal{L}$ in Eq. (22) as summarized in Proposition 1 in the main text. Next, we show how we can use the methodology proposed in [37] to derive the same ELBO as in Eq. (26).
The joint distribution of the proposed ETGPSSM is given by:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}).$ (52)
The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
where the transition term is specified as:
$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}={\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},\ \mathbf{Q}\right),$ (54)
with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$.
We can employ the following variational distribution to infer the latent variables $\{{\mathbf{x}}_{0:T},\mathbf{w}\}$ and the GP $\tilde{f}$:
$q({\mathbf{x}}_{0:T},\tilde{f},\mathbf{w})=q(\tilde{f})q(\mathbf{w})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1}),$ (55)
which mirrors the generative structure of the model in Eq. (52) and assumes independence between $\tilde{f}$ and $\mathbf{w}$.
Given the joint distribution of our model in Eq. (52) and the variational distribution in Eq. (55), augmented by the sparse GP, we can formulate the ELBO $\mathcal{L}_{2}$ according to the general definition provided in Eq. (51). The detailed derivations are presented as follows.
$\mathcal{L}_{2}$ $=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f},{\mathbf{u}})}{q({\mathbf{x}}_{0:T},\mathbf{w},\tilde{f},{\mathbf{u}})}\right]$ (56a)
@@
$=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (56g)
which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [37]:
Assumption 1.
$q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t}).$ (57)
Assumption 2.
$p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}).$ (58)
These two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP. The rationale behind these approximations is as follows:
Remark 5.
It can be observed that, by adopting the approach from [37], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text. However, our contribution in this paper lies in providing an alternative perspective—one that is more straightforward and assumption-free—by leveraging only the EnKF for inference. This not only simplifies the derivation but also enhances the practicality of the method.